OpenStudy (anonymous):

Last year Tammy had $4,500 to invest in two savings accounts. One account earns 7% simple annual interest, and the other account earns 6% simple annual interest. At the end of the year, Tammy had earned a total of $290 in interest from the two savings accounts. Part A Write a system of two equations that could be used to find the amount of money invested in each account. Let x represent the amount invested at 7%, and let y represent the amount invested at 6%. Part B Solve the system of equations. How much did Tammy have invested at each rate?

OpenStudy (neer2890):

tammy had total 4500 to invest. so if he invests $x for 7% interest and $y for 6%. then x+y=4500................(i) also, 0.7x+0.6y=290.................(ii) solve the equations to get x and y.

OpenStudy (anonymous):

x=-24100 and y=28600

OpenStudy (neer2890):

sorry. 2nd equation should be 0.07x+0.06y=290

OpenStudy (anonymous):

oh so then it would be x=2000 and y=2500

OpenStudy (neer2890):

yeah..! correct

OpenStudy (anonymous):

How would I write out part 2?

OpenStudy (neer2890):

we have already solved part B so you can write that tammy have invested $2000 at 7% interest per annum and $2500 at 6% interest per annum.

OpenStudy (anonymous):

ohhhh okay! Wow thank you! I learned a lot off of this (:

OpenStudy (neer2890):

you're welcome....:)